Your investments increased in value by 12.3 percent last year but your purchasing power increased by only 8.3 percent. What was the approximate inflation rate? (Enter your answer as a percent rounded to 1 decimal place. Omit the "%" sign in your response.)

      

  Inflation rate %
Answer #1

Here,

Nominal Rate of Return = 12.30%

Real Rate of Return = 8.30%

(1 + Nominal Rate) = (1 + Real Rate)(1 + Inflation Rate)

(1 + 0.123) = (1 + 0.083)(1 + Inflation Rate)

Inflation Rate = (1.123/1.083) - 1

Inflation Rate = 3.7%